OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1 1/5 cups of juice. How many cups of juice are there per serving? A. 6/5 B. 1/5 C. 1/6 D. 5/6

OpenStudy (anonymous):

i need help ali

OpenStudy (anonymous):

sorry cant help yoou now i love you bye! :p

OpenStudy (anonymous):

hallo

OpenStudy (anonymous):

i can see ur halo

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1 1/5 = 6/5 6/5 divided by 6 (that's how many servings there are) = 1/5

OpenStudy (anonymous):

lol thx

OpenStudy (anonymous):

i have a other one

OpenStudy (anonymous):

A recipe that makes 8 servings calls for cup of flour. Jeff modifies the recipe so that it can serve 10. How many cups of flour does Jeff need? A. B. C. D.